OpenStudy (anonymous):

HELP. Given that IJ=x-4, IK=9x-9, and JK=5x+ 12. Find IJ,IK,andJK.

OpenStudy (anonymous):

 so (x-4)+(5x+12)= 9x-9

OpenStudy (anonymous):

then 6x-8=9x-9 1/3=x then plug in 1/3 into each equation where the x is

OpenStudy (anonymous):

@Lexiskm do you understand?

OpenStudy (anonymous):

wait so we just plug 1/3 in?

OpenStudy (anonymous):

yes